The Linux Distribution HOWTO

maintained by Eric S. Raymond, esr@snark.thyrsus.com v4.1, 7 June 1995

This document lists the various software distributions and services available for Linux via mail order, anonymous FTP, and other sources. You are looking at version 4.1, $Date: 1995/06/07 14:09:53 $

Contents

1 Introduction 2 1.1 Document History and the Editorial We . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2 New versions of this document . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3 And so on... . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2 General Information About Linux Software Distributions 4

3 Network Distributions 4 3.1 Slackware Linux Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.2 MCC Interim Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.3 TAMU Linux Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.4 The BOGUS Linux Release . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.5 Debian Linux Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

4 Commercial Linux Distributions in the US 9 4.1 Yggdrasil Plug-and-Play Linux CD-ROM and the Linux Bible . . . . . . . . . . . . . . . . . . 9 4.2 Professional Slackware CD-ROM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 4.3 Linux from Nascent CD-ROM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 4.4 Red Hat Software Linux CD-ROM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

5 European Distributions 14 5.1 Linux Support Team Erlangen Distribution (LST) . . . . . . . . . . . . . . . . . . . . . . . . 14 5.2 S.u.S.E. GmbH German Linux CD-ROM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 5.3 Unifix 1.02 CD-ROM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 6 Re-Packagers 18

6.1 InfoMagic Developer’s Resource CD-ROM kit . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 6.2 Morse Linux Quarterly CD-ROM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 6.3 Trans-Ameritech Linux plus BSD CD-ROM . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 6.4 Linux Systems Labs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

7 Resellers 23 7.1 Spheric Microsystems Discount Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 7.2 Universal CD-ROM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 7.3 Lasermoon Ltd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 7.4 Takelap Systems Ltd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

8 Miscellany 26 8.1 Unifix Custom Linux CD-ROM Mastering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 8.2 Linux Journal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 8.3 Sequoia International Motif Development Package . . . . . . . . . . . . . . . . . . . . . . . . 27 8.4 SSC Books and Linux Documentation Project Manuals . . . . . . . . . . . . . . . . . . . . . 28 8.5 Extent Verlag, LDP Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 8.6 UPython SLS and Slackware Diskette Labels . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

9 Submissions To This Document 30 9.1 Types of submissions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 9.2 How to submit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 9.3 Submission Guidelines¿ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

1 Introduction

There is no single distribution of the Linux software. Instead, there are many such distributions, available both via anonymous FTP and mail order on floppy, tape, or CD-ROM. This document is an attempt to present information on many of the available distributions of Linux. This document also contains information on many miscellaneous services and goods available for Linux, ranging from consulting and support to Tshirts and virtual beer. The purpose of this document is to provide short summaries of the many Linux distributions and mail-order services, and to provide pointers for the reader to find more information. The information presented here is far from complete; there are many more Linux distributions and services than are listed here. Unfortunately, we have not received submissions from many of the organizations providing these services. Please see 9 near the end of this document for information on making a submission. It’s easy and will take less then five minutes.

1.1 Document History and the Editorial We

This document was originated by Bill Riemers. Matt Welsh maintained the second version. Erik Troan (ewt) maintained the document to release 3.0.

In January 1995, Eric Raymond, while unaware of the existence of this document, began to develop a similar FAQ in the format of his now-discontinued ”PC-clone UNIX Software Buyer’s Guide”, which had covered mainly System V UNIXes and BSD/OS.

In March 1995, Eric approached Erik about cooperating on a merged version. In early April, Erik went to work for Red Hat Software, and (wishing to avoid a conflict of interest) handed the document to Eric. Eric merged in a lot of new information and added several new fields to the distribution entries.

Accordingly, this document is a sort of serial collaboration. The editorial ‘we’ generally tags observations by all the maintainers; ‘I’ is Eric (the current one) speaking.

We are delighted to acknowledge the contributions of all the Linux users and Internet hackers who have contributed information and feedback.

1.2 New versions of this document

New versions of this document may be posted periodically to the newsgroups comp.os.linux.answers and comp.os.linux.misc. The document is archived on a number of Linux FTP sites, including

sunsite.unc.edu in pub/Linux/docs/HOWTO.

You can also view the latest version of this on the World Wide Web via the URL

(http://sunsite.unc.edu/mdw/HOWTO/Distribution-HOWTO.html).

1.3 And so on...

Disclaimer: We make absolutely no guarantee as to the correctness of the information, prices, and ordering details given in this document. Check the last modified field of each to get an idea of its currency. Furthermore, unless otherwise stated the Linux software comes with ABSOLUTELY NO WARRANTY .

We do not endorse or work for any of the distributors listed in this document. We merely include their information here as a service to the Linux community, as a “buyer’s guide”. Inclusion in this document does not mean that we recommend any of the products listed here.

Feel free to mail any questions or comments about this HOWTO to Eric S. Raymond, esr@snark.thyrsus.com. Please do not send me general Linux questions or requests for help in choosing a distribution unless you’re willing to hire me at normal consulting rates; I don’t have time to deal with them, and I try to put everything I know about choosing a distribution in this document.

And now for the good stuff.

2 General Information About Linux Software Distributions

The current de facto standard distribution is the Slackware release, available freely over the net. Slackware is easy to install and fairly complete. It’s also generally up-to-date with current versions of Linux software. The Linux Installation HOWTO documents installation of Slackware.

If you have a CD-ROM drive, you have many more distributions to choose from. We haven’t personally reviewed most of them, and you’re more likely to get help from people on the Net for network-available distributions such as Slackware. On the other hand, some CD-ROM distributions come with documentation and support of their own.

Some distributions, such as Slackware Pro, Red Hat Linux, and Yggdrasil take advantage of the storage space available on CD-ROM to minimize the amount of hard disk space Linux takes.

3 Network Distributions

In this section, we cover distributions available over the network for free. Many of the CD-ROM are simply re-packaged versions of net distributions.

3.1 Slackware Linux Distribution

Distributor:

Patrick Volkerding

     Net: volkerdi@mhd1.moorhead.msus.edu (mail:volkerdi@mhd1.moorhead.msus.edu)

Provider’s Description:

Slackware Linux is a full featured distribution of the Linux operating system designed for 386/486 computers with a 3.5" floppy. Slackware changes rapidly, but here’s a current (partial) feature list:

A:

Base Linux system

AP:

Various applications that do not need X

D:

Program Development (C, C++, Kernel source, Lisp, Perl, etc.)

E:

GNU Emacs (19.27, with and without X11 support. Your choice.)

F:

FAQ lists (last updated: Oct 1994)

I:

Info pages. IV:

Interviews: libraries, include files, Doc and Idraw apps for X N:

Networking (TCP/IP, UUCP, Mail) OOP:

Object Oriented Programming (GNU Smalltalk 1.1.1) and X11 interface Q:

All kinds of extra kernels (based on Linux 1.1.59) T:

TeX TCL:

Tcl/Tk/TclX, Tcl language, and Tk toolkit for developing X apps X:

XFree-86 3.1 Base X Window System XAP:

Applications for X XD:

XFree-86 3.1 X server development, PEX extensions, and man pages XV:

XView3.2p1-X11R6. (OpenLook virtual Window Manager, apps) Y:

Games (that do not require X) Internet Access:

The home site is ftp.cdrom.com, where the latest distribution can be found in /pub/linux/slackware. Ordering:

FTP only, although various independent distributors provide it on disk, floppy, and CD. Entry last modified:

14 Dec 1994 Editor’s Comments:

A lot of commercial Linuxes are re-packagings of Slackware (see the Re-Packagers section below.

3.2 MCC Interim Linux

Distributor:

Dr. A. V. Le Blanc University of Manchester

Net: LeBlanc@mcc.ac.uk (mailto:LeBlanc@mcc.ac.uk)
FTP: (ftp://ftp.mcc.ac.uk)

Provider’s Description:

Base Linux installation. Complete sources and patches for all included software are available.

Full details vary from release to release; see /pub/linux/mcc-interim/*/Acknowledgements at

ftp.mcc.ac.uk. Roughly: this includes the kernel (with source), C, C++, groff, man pages, basic utilities, networking. The binaries fit on about 10 floppies. It is also possible to have them on a DOS or Linux partition on a hard drive (plus one floppy) or on an NFS-mountable partition (plus two floppies).

Internet Access:

By anonymous ftp from ftp.mcc.ac.uk; mirrored at tsx-11.mit.edu, nic.funet.fi,
sunsite.unc.edu, and elsewhere. At ftp.mcc.ac.uk, in /pub/linux/mcc-interim.

Miscellaneous:

Suggestions and contributions are welcome.

Last Freeze Date:

Unknown. New releases approximately every six months.

Entry last modified:

May 15 1995.

Editor’s comments:

The current MCC-Interim release is based on Linux 1.0 with a few patches. If you are one of those people who demands to be on top of the current release, this release isn’t for you. However, if you’re looking for a small and stable Linux system, this makes a good starting point. Dr. LeBlanc adds: ”And if you do demand to be on top of the current release, you can get there from MCC Interim quite easily. For example, to upgrade to the current kernel you need only the kernel sources, bdflush, and a new version of procps. Of course, you will probably want other things as well... One of the big disadvantages of MCC is that it doesn’t include X or packages that many people want, such as perl or tck/tk. Strictly speaking, the distribution has always been intended for internal use at MCC, and the fact that many people find it useful is serendipitous. Less strictly speaking, I would say that it works better for developers than for people who know nothing about Unix. I recommend the simpler commercial distributions and Bogus to people who know nothing.

3.3 TAMU Linux Distribution

Distributor:

Dave Safford Texas A&M University,

Net: dave.safford@net.tamu.edu (mailto:dave.safford@net.tamu.edu)

Provider’s Description:

TAMU.99p12+ is the latest release in the TAMU linux series. Unlike previous releases, this one includes both integrated source and binary sets, with the entire binary set created from a single top level source make. This ensures that all programs are compiled and linked with the same current tools and libraries, and guarantees availability of working source for every program in the binary set. In addition, the new boot diskette fully automates the installation process, including partitioning, lilo bootstrapping, and network configuration. Installation requires no rebooting, and requires the user to know only the host’s name and IP address. At every step of installation, the program provides intelligent defaults, making it a snap for novices, while allowing experts full flexibility in setting installation parameters. Reliability has been improved over past TAMU installation by the use of labels on all disk images, so that the program can detect and recover from bad or misordered diskettes. This release is a full-featured package, including XFree86-1.3, emacs-19.18, net-2, bootutils, and sources for all installation programs (without any use restrictions :-).

Internet Access:

The latest TAMU release is available by anonymous ftp from net.tamu.edu:pub/linux.

Entry last modified:

31 Oct 1993.

3.4 The BOGUS Linux Release

Distributor:

The BOGUS Group

Net: linux-bogus@cs.unc.edu (mailto:linux-bogus@cs.unc.edu)

Provider’s Description:

BOGUS is an unsupported, ftp-able distribution targeted at competent Linux hackers who want a stable development system. If you do not have experience developing, installing, and maintaining a Linux system, BOGUS is probably not for you. BOGUS is a ”bootstrapping” release: all of the patches needed to compile the complete system are provided, including explicit instructions on how we did the compilation. Indeed, the final build is totally automatic, and required the user to invoke a single command.

BOGUS has many other features, but the inclusion of pristine source, detailed patches, and a utility to support automatic patch application and compilation is the most important feature of BOGUS. For more information and a complete list of features, license restrictions, and other details, please see the Announce-BOGUS-* files at the ftp sites.

Internet Access:

The following sites are official mirror sites for the BOGUS release:

• ftp.cdrom.com:/pub/bogus

• ftp.gwdg.de:/pub/linux/bogus

• ftp.cps.cmich.edu:/pub/linux/packages/bogus

• ftp.nvg.unit.no:/pub/linux/bogus

• ftp.thepoint.com:/pub/linux/BOGUS

• ftp.ibp.fr:/pub/linux/distributions/bogus

• sunsite.unc.edu:/pub/Linux/distributions/bogus

Ordering:

BOGUS 1.0.1 is also available on the Trans-Ameritech Systems ”Linux Plus BSD CD-ROM” (Fall Supplement 2, November 1994) for $19.95. Trans-Ameritech may be contacted at: 408.727.3883 (voice), 408.727.3882 (fax), or order@trans-am.com (mailto:order@trans-am.com).

Miscellaneous:

To join the BOGUS mailing list, send mail to majordomo@cs.unc.edu (mailto:majordomo@cs.unc.edu) with ”subscribe linuxbogus” in the body of the message (i.e., not in the subject). When you do this, make sure that you have a valid return address – many people who set up a mailer for the first time have a return address which looks like joe@pizza instead of joe@pizza.cs.unc.edu. If you interact with the majordomo server successfully, you will receive mail. If you don’t receive mail in about an hour, please don’t continue to use the majordomo server as test bed. Please note that the BOGUS release is COMPLETELY UNSUPPORTED: mail sent to the mailing list may not be answered, bugs may not be fixed, and it may be quite a while before there is another BOGUS release.

Entry last modified:

14 Dec 1994.

3.5 Debian Linux Distribution

Distributor:

Ian A. Murdock The Debian Linux Association Station 11 P.O. Box 3121 West Lafayette, IN 47906, USA

Net: imurdock@gnu.ai.mit.edu (mailto:imurdock@gnu.ai.mit.edu)

Provider’s Description:

Debian Linux is an effort to create a well thought-out, powerful, flexible yet complete Linux distribution. The motivations behind Debian are detailed in the Debian Manifesto, a copy of which can be acquired as described below. For up-to-date information, please look at the files in the directory

/pub/Linux/distributions/debian/info at sunsite.unc.edu. If you do not have access to the

Internet or FTP, you may obtain printed copies by sending a self-addressed, stamped envelope to the contact address given above. This information includes how Debian can be obtained (via FTP or mail-order), why Debian is being constructed (the Manifesto) and other general information (including how to join the Debian mailing lists).

Internet Access:

Debian has not yet been ”officially” released, but BETA releases are available to the general pub-

lic at sunsite.unc.edu in the directory /pub/Linux/distributions/debian. It is mirrored at
ftp.mcc.ac.uk.

Entry last modified:

21 Feb 1994.

Editor’s comments:

The Debian distribution is currently under development by a “semi-closed” group of developers (rather then an individual) and has been adopted as the Free Software Foundation’s Linux distribution.

Network feedback Dr. A.V. LeBlanc, the priginator of MCC Interim Linux, comments: ”BOGUS is a particularly welldesigned release, and one which I feel is not sufficiently appreciated.”

4 Commercial Linux Distributions in the US

These distributions are available on CD-ROM. They are custom-assembled, rather than just being repackagings of Slackware or some other network release.

4.1 Yggdrasil Plug-and-Play Linux CD-ROM and the Linux Bible

Distributor:

Yggdrasil Computing, Incorporated 4880 Stevens Creek Blvd., Suite 205 San Jose, CA 95129-1034

Net: info@yggdrasil.com (mailto:info@yggdrasil.com)
WWW: (http://www.yggdrasil.com)

Vox: (800) 261-6630, (408) 261-6630 Fax: (408) 261-6631

Provider’s Description:

Yggdrasil Plug-and-Play Linux is a complete CD-ROM distribution of the Linux operating system. It includes a great deal of software—nearly every package that you would expect to find on a complete UNIX system is available. A complete file list is available via FTP from yggdrasil.com. The Linux Bible is a compendium of Linux documentation, including 5 books from the Linux Documentation Project, The Yggdrasil installation manual, and the complete set of Linux HOWTO guides. We publish this and offer this separately from the CD-ROM distribution. At the top of the sophistication hierarchy, programmers who want to explore or add the occasional feature know that Yggdrasil is the Linux distribution with a fully buildable source tree and with the ability to automatically trace installed files back to their sources. Everybody, especially new users, will appreciate the Plug-and-Play operation for which Yggdrasil’s product is named. Put the media in a computer with supported hardware, turn the computer on, and it’s running everything, straight from the CDROM. (This is no longer unique you Yggdrasil, but Yggdrasil did it first.) The login screen lists a number of preconfigured user names, including ”install”, which installs the system, giving paragraphs of explanation about every question that it asks the user. X windows configuration is automated too, prompting the user for configuration information the first time xinit is run. From X windows, a graphical control panel allows simple “fill in the blanks” configuration of networking, SLIP, outgoing UUCP, the printer, NNTP, and many other features that previously required the knowledge of a system administrator to configure. The next release will feature ELF object-code format. Yggdrasil has been working hard on improving the installation (making packages smaller, etc.). For more information, send mail to info@yggdrasil.com, surf to www.yggdrasil.com, FTP to yggdrasil.com, or contact us by any convenient method.

Ordering:

Plug-and-Play Linux costs $34.95 is and available directly from Yggdrasil or from your local computer, software or technical book store. If Plug-and-Play Linux is not available from your favorite reseller, help promote Linux by making it your mission to change that. Give your reseller our phone number and demand that they carry Plug-and-Play Linux.

Yggdrasil offers a $7 discount for upgrades or crossgrades. Send us your old Yggdrasil release and a check for $27.95 + $5 shipping and handling to upgrade. Or do the same, but send us a competing distribution such any version of SCO, Esix, minix, or one of the CD’s with the slackware floppy images, and tell us where you got it, so that we can make sure that your favorite reseller carries our products too.

Miscellaneous:

Yggdrasil also sells OSF/Motif and The Linux Bible, a compilation of works from the Linux Documentation project. In addition to a copy of the Plug-and-Play Linux manual, The Linux Bible includes Linux Installation and Getting Started, Network Administrator’s Guide, Kernel Hacker’s Guide, System Administrator’s Guide, Linux User’s Guide, and HOWTO guides on hardware compatibility, distributions, dos emulation, ethernet, floppy tape, installation, mail, networking version 2, MGR (an alternative to X windows), X windows, usenet news, printing, SCSI, serial communications, sound, and UUCP. The Linux Bible costs $39.95 and is printed on recycled paper. $1 from every copy sold is donated to the Linux Documentation Project. OSF/Motif costs $149.95, $5 of which is donated to the development of a free Motif clone.

Last Freeze Date:

December 1 1994. Another release is scheduled for late May.

Entry last modified:

27 April 1995

Editor’s Comments:

Yggdrasil has a mixed reputation. On the one hand, they include lots of well-chosen stuff with their distribution; the CD-ROM-only demo mode is impressive; installation and administration are easy enough to live up to their ”Plug And Play” name. The 95-page manual is very nice. On the other hand, the installation script leaves a lot of shambolic links lying around; many applications (including some as basic as file(1)) won’t run without the CD-ROM mounted. And don’t even try a full installation with less than a 1GB root disk.

4.2 Professional Slackware CD-ROM

Distributor:

Morse Telecommunications, Inc. 162 West Park Avenue, 2nd floor Long Beach, NY 11561

Net: order@morse.net (mailto:order@morse.net) (orders)

support@morse.net (mailto:support@morse.net) (info)
WWW: (http://www.morse.net)
FTP: (ftp://www.morse.net)

Vox: (800) 60-MORSE (orders), (516) 889-8610 (tech support)

Fax: (516) 889-8665

Hours: 9AM - 5PM EST, Monday through Friday. Fax available 24 hours.

Provider’s Description:

Slackware Professional Linux 2.2.0.1 is a 4 CD-ROM set (shipping early May). This set also includes the contents of the sunsite.unc.edu archives and a 600-page printed manual. Thirty days of technical support (clock starting with the first phone call) is provided with this set at no additional charge.

Ordering:

Orders may be placed via phone, fax or email as listed above. The Professional version is $68.95 plud $5 per order for shipping and handling. All orders generally ship the same day if received by 4PM EST. A full product description is available by email or at the Web site. See Morse’s entry under Re-Packafers for the Morse Quarterly CD-ROM.

Last Freeze Date:

April 15 1995. Future releases will track Slackware.

Entry last modified:

27 April 1995

4.3 Linux from Nascent CD-ROM

Distributor:

Nascent Technology Linux from Nascent CDROM P.O. Box 60669 Sunnyvale CA 94088-0669 USA

Net: nascent@netcom.com (mailto:nascent@netcom.com)

Vox: (408) 737-9500 Fax: (408) 241-9390

Provider’s Description:

The Linux from Nascent CDROM is a new distribution of the Linux operating system which includes over 400 mbytes of source code, binaries, and documentation for Linux and applications. It features automated root, swap, package, network, and user account installation from CDROM. Linux can be can be run directly from the CDROM and floppy. The Nascent CDROM features Xwindows, Openlook, TeX, GNU compiler and utilities, Magic and Spice electronic design tools, and over 100 high resolution images translated from Kodak PhotoCD(TM). Each source archive is distributed with an associated notes file to allow you to browse and install applications using a consistent interface.

A listing of the contents of the Nascent CDROM as well as a current copy of the CDROM announcement and order form may be obtained via anonymous ftp at netcom.com:/pub/na/nascent. New features in the 2.0 release:

• SysV initialization including runlevel support

• Conformance to the Linux Filesystem Standard FSSTND

• Hierarchical source build

• Automated package installation from CDROM, harddisk, or network

• Laptop Card Services including modem and networking support

• Support for OSF/Motif3d window manager (provided separately)

• Electronic design tools for behavioral synthesis and sea-of-gates place and route

• Andrew and Tcl/Tk

• UUCP, mail and news Ordering:

The Linux from Nascent CDROM, Version 2.0, is only $59.95 plus shipping and handling. Nascent also offers the Linux from Nascent Plus package for only $119.95, which includes six months of email support and a 30 future release of the CDROM with your CDROM purchase. Nascent accepts Mastercard, VISA, checks, and money orders as payment. To order your Linux from Nascent CDROM, mail, email or fax a completed order form to Nascent.

Last Freeze Date:

Unknown. Entry last modified:

27 April 1995 Editor’s Comments:

Nascent’s ads indicate both visually and textually that this Linux is targeted to hardware and IC designers.

4.4 Red Hat Software Linux CD-ROM

Distributor:

Red Hat Software PO Box 4325. Chapel Hill NC 27515

Net: redhat@redhat.com (mailto:redhat@redhat.com)
WWW: (http://www.redhat.com)

Vox: (919)-309-9560 Provider’s Description:

RHS Linux contains over 650Mb of binaries and source code and comes with a manual that covers the installation process and the RHS Linux system administration tools. In addition the manual contains the entire Linux Installation and Getting Started book by Matt Welsh, and Ian Jackson’s Linux FAQ. Included with the purchase of RHS Linux is 30 days of free installation support, and a card for a complimentary copy of the Linux Journal. More details are available from our FTP site.

Ordering:

Available directly from Red Hat Software and most distributors of Linux related products. Please contact Red Hat Software for ordering details. (see contact info above). Suggested retail price is $49.95.

Last Freeze Date:

Unknown.

Entry last modified:

27 April 1995.

Editor’s Comments:

If you call Red Hat (as of late April) you’ll get a message informing you that they’re out of stock. They inform us that a new version is expected out about May 15 1995.

5 European Distributions

These are distributions intended for use by non-English-speakers. So far, all happen to be German.

5.1 Linux Support Team Erlangen Distribution (LST)

Distributor:

The Linux Support Team Erlangen (a small group of students at the University of Erlangen-Nuernberg). Contacts:

Stefan Probst (snprobst@cip.informatik.uni-erlangen.de) or
Ralf Flaxa (rfflaxa@informatik.uni-erlangen.de).

Probably, there will be a collective address soon.

Provider’s Description:

The LST distribution’s goal is to provide a solid, reliable, easy to install (even for beginners) and well-documented system. We are not hunting for the newest kernel or gcc versions. We do updates when they are necessary or provide really new functionality, are well tested, integrated in the system and working smoothly with the rest of the system. The distribution consists of a base system and additional packages. Currently the following packages are available: doku (doc), text, tex, develop, xdevelop, xbasis, xappl, xemacs, tinyx, network, grafik, src, misc.

The complete system is 50 HD disks and 1500 pages printed documentation including the LDP guides (IGS,KHG,NAG), HOWTOs, FAQs, the German Linuxhandbuch, install-guide and many other useful documents we collected over the time. Our distribution is preconfigured for German users and comes with a 50 page step by step installation guide that leads you through the menu-driven installation scripts (both in German). We started this distribution to help newcomers with their first steps into Linux. Therefore our scripts are smart enough to handle all of the “dirty work” like setting up system configuraton (including LILO, modem, mouse, mounts, mtools, access to DOS, users, X11) and network configuration (TCP/IP, Routing, Mail, News, UUCP, SLIP). Historically (and of course easier for us) most of the documentation for this distribution is written in German, sorry, but we are planning to translate those documents in english, but at the moment we don’t have time for this. Any volunteers are welcome! Our scripts also accept to install SLS and Slackware packages, but with no warranty if they work well together with the rest of the system.

Internet Access:

Via anonymous ftp: ftp.uni-erlangen.de under pub/Linux/LST.Distribution, or on 3.5" disks

(also on one QIC-80 tape).

Ordering:

Order at following adress: Stefan Probst, In der Reuth 200, 91056 Erlangen, Germany.

• complete documentation (IGS, LHB, KHG, GDB, NAG, HOWTOs, FAQs, Installguide) together over 1500 pages for DM 139,– (plus shipping).

• complete documentation plus complete distribution on QIC-80 tape (about 65 MB) for DM 199,– (including shipping in Germany).

• complete documentation plus complete distribution on about 50 3.5" disks for DM 269,– (including shipping in Germany).

Entry last modified:

21 Dec 1993.

5.2 S.u.S.E. GmbH German Linux CD-ROM

Distributor:

S.u.S.E. GmbH Gebhardtstrasse 2 90762 Fuerth Germany

Net: bestellung@suse.de (mailto:bestellung@suse.de)

Vox: +49 911-74053-31 Fax: +49 911-7417755 Provider’s Description:

This is a CD especially for German users. It contains a German version of the Slackware Distribution and two additional german distributions based on Slackware and SLS. It contains additional software, such as the Postgres system, the pbm-Tools and lots of other tools and utilities (ctwm window manager, GREAT environment, andrew toolkit, the POV-ray raytracer and so on). The source code of the complete system is included, too, as is a live system to enable using software without the need to install everything to the harddrive. The current version of the CD contains version 1.2.0 of the Slackware distribution containing version 1.0 of the linux kernel, version 2.5.8 of the GNU-Compiler and XFree86tm 2.1. A German installation guide will help new users to successfully install the system. The CD is updated every three months, so the next one will be released in July 1994.

Ordering:

By mail, see contact information above. Price: The price for a single CD is 89,– DM for new customers. If you are already a S.u.S.E.-customer, you will get the CD for an update-price of 78,– DM. There is also the posibility to subscribe for one year; then you will receive 4 CDs for the price of 230,– DM. Disks: Of course, you can get the German version of the Slackware Distribution on 3,5" disks. You can get either a basis version of the distribution (about 28 disks) for 89,– DM, the ”standard” version (about 38 disks) for 119,– DM or the ”full” version (about 59 disks) for 159,– DM. The update prices are 69,– DM, 94,– DM and 139,– DM respectively. Documentation: S.u.S.E. is selling high-quality printed issues of the books from the Linux Documentation project (LDP) for people that would like to have a real book and not only printed loose papers. The prices are 24,– DM for Matt Welsh’s Linux Installation and Getting Started, 39,– DM for Olaf Kirch’s Network Administrators’ Guide and 29,– DM for Michael K. Johnson’s Kernel Hackers’ guide. Additionally, S.u.S.E. offers the complete O’Reilly series which is the best documentation available for Unix in general. Of course you can order the well known German Linux books: Deutsches Anwenderhandbuch for 49,– DM and Linux - vom PC zur Workstation for 38,– DM. Motif: Metrolink Motif for Linux is available for the price of 169,– DM for the runtime system only and for the price of 288,– DM you will get the runtime and development system.

Support:

Service and support for our customers by Email, Fax, snail mail and phone (hotline Monday and Thursday from 13:00h-17:00h.)

Miscellaneous:

Free info material and catalogues of CDROMs and books. (All Infomagic Unix CDs available). Selling commercial software for Linux. Development of individual software for Linux and other Unix systems. Almost anything concerning Linux - just ask!

Entry last modified:

23 May 1994

5.3 Unifix 1.02 CD-ROM

Distributor:

Unifix Software GmbH Postfach 4918 D-38039 Braunschweig, Germany

Vox: +49 (0)531 515161 Fax: +49 (0)531 515162

Provider’s Description:

This is a Linux CD-ROM distribution with emphasis on easy and fast installation. Though it is possible to install everything on the hard disk, we do not recommend doing so, because it runs fast enough directly from CD. For example starting emacs the first time (in text mode) takes:

                From single speed Mitsumi: 24 s
                From double speed Mitsumi: 11 s
                From Toshiba 3401:          7 s

The dynamic buffer cache of Linux takes care of the second and subsequent times a program is started, which needs less than one second. Our system requires about 5 MB on the hard disk for configuration files and system management tools. Additionally we recommend a 16 MB swap partition. The distribution contains most of the standard programs in current versions, e.g. Linux 0.99.15e, XFree 2.0, emacs 19.22 and gcc 2.5.7. Full preconfigured source for everything is included. Because Unifix is a European distribution it provides full support for iso8859-1 character sets. From the shells and emacs through ls, TeX and the print system everything supports 8-bit characters. Printing is supported through System V compatible printer drivers which can be controlled through lp’s -o options. Printer drivers are included for dumb text mode printers, for postscript- or ghostscriptcompatible printers and for networked printing. These drivers know how to guess the type of the files so e.g. compressed man pages or DVI files can be printed directly.

Ordering:

Unifix is available only on CD. It comes in a Unifix/Linux binder with two boot floppies and about 70 pages of installation instructions in german. An english version will be available Apr 94. The price is 159 DM (about $100) and includes taxes and shipping; Eurocard/Mastercard/Visa are accepted.

Entry last modified:

18 Feb 1994.

6 Re-Packagers

This section lists re-packagers of the Linux distributions listed above. In other words, the people selling the software below probably do not maintain or support the software itself.

6.1 InfoMagic Developer’s Resource CD-ROM kit

Distributor:

InfoMagic, Inc. PO Box 30370. Flagstaff, AZ 86003-0370

Net: info@infomagic.com (mailto:info@infomagic.com)
WWW: (http://www.infomagic.com)

Vox: (800)-800-6613 (orders), (520)-526-9852 (tech support) Fax: (520)-526-9573

Provider’s Description:

The InfoMagic Linux Developer’s Resource is a complete snapshot of the sunsite.unc.edu and tsx-11.mit.edu archives. It also includes the complete GNU software collection (in source form). The following Linux distributions are included on the discs: Slackware, Bogus, Debian, SLS, TAMU, MCC, and JE (Japanese Extensions). The Slackware distribution has been completely unpacked allowing many packages to be run directly from the disc. Sources for all the packages in Slackware are also included. The Linux HOWTO documents have been formatted for use with the Microsoft Multimedia Viewer (which is included) to allow browsing and full-text search under Microsoft Windows. InfoMagic also includes complete archives of the linux mailing lists. InfoMagic also distributes the ”Installation and Getting Started Guide”, by Matt Welsh as well as Linux T-Shirts.

Ordering:

The 3-CD set is $25/copy. Shipping within the US is $5 (USPS Priority Mail), outside the US $10 (International Airmail), FedEx and UPS on request. Orders may be placed via phone, fax, or email (a PGP key is available: finger orders@InfoMagic.com). We accept Visa, Mastercard, and AMEX. 1year subscriptions are available for $150 (within the US) and $175 (outside the US) including shipping. A 1 year subscription consists of 6 releases, one every two months or so.

Internet Access:

The contents of the CD’s may be found at either (http://www.infomagic.com),

(ftp://ftp.uu.net/vendor/InfoMagic/cd-roms/linux).

Support:

We provide installation tech support by phone and email. The set includes a 30-page quick-start. No support contracts yet.

Last Freeze Date:

21 Mar 1995. Updates about every two months. The Slackware release is 2.2.0.

Entry last modified:

27 April 1995.

6.2 Morse Linux Quarterly CD-ROM

Distributor:

Morse Telecommunications, Inc. 162 West Park Avenue, 2nd floor Long Beach, NY 11561

Net: order@morse.net (mailto:order@morse.net) (orders)

support@morse.net (mailto:support@morse.net) (info)
WWW: (http://www.morse.net)
FTP: (ftp://www.morse.net)

Vox: (800) 60-MORSE (orders), (516) 889-8610 (tech support) Fax: (516) 889-8665

Hours: 9AM - 5PM EST, Monday through Friday. Fax available 24 hours.

Provider’s Description:

The Linux Quarterly CDROM contains the complete contents of tsx-11.mit.edu, one of the most popular Internet Linux sites which provides both source and binaries of major Linux distributions, utilities, source code and documentation. The CD-ROM also includes the complete contents of prep.ai.mit.edu, the repository of the FSF’s GNU source archives, is included. For first-time users, the Quarterly CDROM contains a Microsoft Windows(tm) front-end to assist the user in creating boot disks that can be used to install Linux directly from The Linux Quarterly CDROM. Support for the UMSDOS filesystem is now also included. This gives the user the ability to install Linux directly to an MSDOS system without having to repartion their hard drive. We also sell a Slackware Professional Linux 2.2 which is a 4 CD-ROM set (shipping early May). See above under Commercial Distributions.

Ordering:

Orders may be placed via phone, fax or email as listed above. The Linux Quarterly CDROM is priced at $29.95 plus $5 shipping and handling per order. Existing customers may upgrade for $22.95 plus S/H.

Miscellaneous:

The contents of the Quarterly CDROM, the cover artwork in JPEG format, and the text of the USENET announcement can be found on the following FTP sites:

             tsx-11.mit.edu:/pub/linux/advertisements/TLQ-Spring94.tar.z
             sunsite.unc.edu:/pub/Linux/distributions/cdrom/TLQ-Spring94.tar.z

Last Freeze Date:

November 1994. Releases every three months.

Entry last modified:

27 April 1995

6.3 Trans-Ameritech Linux plus BSD CD-ROM

Distributor:

Trans-Ameritech 2342A Walsh Avenue Santa Clara, CA 95051

Net: info@trans-am.com (mailto:info@trans-am.com)

WWW: <none> Vox: (408)-727-3883 Fax: (408)-727-3882

Provider’s Description:

Trans-Ameritech has published 6 releases of Linux by the end of 94. CD-ROM based on the Slackware distribution of Linux with all the sources plus full uncompressed filesystem plus NetBSD source and binary distribution. The current full release is Release 4 and a Supplement 2 is also available with updates and additional software. The previous releases cover Linux from 0.99 p.9 (on the SLS release) to the current versions and have FreeBSD, NetBSD and the original 386BSD.

• The installation is highly automated by the Slackware scripts fine tuned to this CDROM but leaves a lot of room for customization. New Linux users will appreciate the DOS based menu driven program that helps to select and prepare the right boot floppy.

• To minimize the possibility of hardware conflicts many extra kernels are provided for different configurations. They are usable for installation and normal operation.

• Many on-line documents are provided for quick reference, including the Linux Documentation Project files in source, dvi and ps formats.

– A lot of applications are included:

– The C/C++ compiler GNU GCC ,

– GNU and international versions of the ispell spell-checker.

– The communications apps: term, minicom, Seyon (X-Windows based).

– Editors: elvis (vi clone), joe, jove

– PostScript clone ghostscript

– Network package with news and email:

– TCP/IP, UUCP, SLIP, CSLIP

– Object oriented GNU Smalltalk, and the Smalltalk Interface to X.

– TCL/Tk (Powerful scripting language with Motif-like X interface)

– Programs for electric engineers and HAM enthusiasts

– Several window managers - openwin, twm, gwm, fvwm.

– Easy X-Windows configuration with many real-life examples.

– Many X applications

– Interviews libraries, include files, and the doc word processor and idraw drawing program.

– Typesetting: TeX, LaTeX, xdvi, dvips, Metafont, groff

– Andrew multimedia word processor with hyperlinks.

– Send and receive fax on either class 1 or class 2 fax modems.

– A lot more.

• To help the first time Linux users many documentation files are provided that are readable from DOS even before installing Linux.

• All the sources are available on the CDROM. The most often needed sources are uncompressed and can be used directly from CDROM.

• An uncompressed Linux filesystem is available for reference and disk space conservation. You can run programs directly from CDROM! There is a large info directory for on-line reference and many manpages.

• For hacker’s reference uncompressed 386 BSD sources are provided.

The latest upgrades and patches are always available through Trans-Ameritech Linux BBS (408)980-9840. It’s free for all Trans-Ameritech customers. Our distribution is targeted for DOS users who want an easy migration path upwards, rather for UNIX experts. We have the only release that can be installed on a DOS-partioned disk without disturbing DOS. After installation, we have a program you can run from COMMAND.COM that boots Linux. In the next release, we will give you the ability to install, configure, and run Linux from within Windows.

Availability and Ordering:

The Internet price for Linux Supplement CD 2 is $20 The price for our current (Release 4) Linux Plus CD is $30 The price for a package order: Linux Plus CD (Release 4) and Linux Supplement 2 CD (Trans-Ameritech Linux Package) in one shipment is - $40 You can order by e-mail, by phone, fax, or by sending a check or money order. If you order with a credit card (VISA, MC, American Express) please indicate the card number, expiration date and your mailing address. The order will be processed and the CD shipped the same day. Shipping and handling in US is $5, Canada/Mexico $6, Overseas $8. COD is available in the US only for $4.50.

California residents please add sales tax. Annual subscriptions (4 releases) are available for $80 plus S&H. (note: there are 4 shipments in a subscription) Example subscription in US is: $80 + $5 x 4 = $100 Subscription in Europe/Japan etc. is: $80 + $8 x 4 = $112

Miscellaneous:

If you have any further questions, please contact us at the above address, or (408) 727-3883, FAX (408)

727-3882, or info@trans-am.com

Last Freeze Date:

Mid-October 1994. Updates about once a quarter.

Entry last modified:

11 Jan 1995.

6.4 Linux Systems Labs

Distributor:

Linux Systems Labs 49884 Miller Ct Chesterfield, MI 48047-2333

Net: info@lsl.com (mailto:info@lsl.com)

WWW: <not yet, up soon> Vox: (810)-716-1700 Fax: (800)-432-0556

Provider’s Description:

We sell Slackware (all 50 disks and I&GS) for $69.95 (updated weekly), resell the Yggdrasil CDROM and bundle the latest version of Slackware with Motif (version1.2.4) for $215.00 (Motif alone for $175.00). We currently print LDP documents duplex on a 600 DPI laser printer. We publish Dr. Linux, an impressive compendium (1170 pages) of the Linux Documentation Project for $49.95 and include a free CD with purchase. A free CD also comes with our other book, Linux: Getting Started (550 pages), which sells for $29.95. Our newest product is Just Logic SQL Database Manager. It is an interactive multiuser SQL server+client for Linux. It also includes C++ class descriptions and API for interfacing with C and C++.

Ordering:

See above for prices. We take MasterCard, Visa and American Express and will discount all products (except SQL) 20% for internet customers.

Last Freeze Date:

8 Feb 1995. New releases monthly. Entry last modified:

9 Feb 1995.

7 Resellers

This section is devoted to distribution houses that re-sell packaged Linuxes.

7.1 Spheric Microsystems Discount Linux

Distributor:

Spheric Microsystems, Inc.

Net: info@spheric.com (mailto:info@spheric.com)

Vox: (800)-869-8649

Provider’s Description:

Spheric has the Linux distribution to fit your tastes, at discount prices. Here’s just a sample, call for latest versions or a catalog:

Slackware Pro 2.0

retail $49.95, only $34.95. (Also available in floppy) Yggdrasil Plug and Play Fall ’94

retail $49.95, only $29.95 InfoMagic Archives

October 1994, kernel 1.1.50, Doom, retail $20.00, only $17.95 Linux Installation and Getting Started Guide

2.1, retail $14.95, only $9.95 Linux Bible - The GNU Testament

retail $49.95, only $39.95 Unix in a Nutshell, O’Reilly reference book

retail $9.95, only $8.95 Unix PowerTools, best tips and tricks

with CD, retail $59.95, only $44.95 The Whole Internet User’s Guide

2nd Edition, retail $24.95, only $19.95 Source Code CDROM

many from source news groups, retail $39.95, only $24.95 C User’s Group CD, volume 100-411

retail $49.95, only $29.95

Ordering:

        Visa, MasterCard, American Express, check, money-orders accepted.
        Call 1-800-869-8649, 9am-9pm Eastern.
        NYS Residents add local sales tax.
        \$5 shipping per order.

Miscellaneous:

Spheric has been distributing Linux since 1992 (back when a few floppies made a great distribution :^). Commercial Linux software, general UNIX books, and CD-ROM subscription plans also available.

Entry last modified:

14 December 1994.

7.2 Universal CD-ROM

Distributor:

Universal CD-ROM 520 Lawrence Express Way, #307 Sunnyvale, Ca. 94086

Net: ucr@bigmall.com (mailto:ucr@bigmall.com)
WWW: (http://www.bigmall.com)

Vox: (408)-992-0453 Fax: (408)-992-0457

Provider’s Description:

We are the publisher of Linux World Book ’95 (World Book ’95 is a complete Linux reference book) and UCR Linux CD. In addition we also carry a complete line of Linux distribution such as: Slackware 2.1, Plug & Play, Walnut Creek CDs, InfoMagic Linux CDs, etc. You can get the latest price list from our WWW site.

Ordering:

We accept: Visa, MasterCard, Money order and Checks. We ship worldwide from Silicon valley.

Entry last modified:

27 April 1995 Editor’s Comments:

The UCR Linux release is a repackaged Slackware.

7.3 Lasermoon Ltd.

Distributor:

Lasermoon Ltd 2a Beaconsfield Road Fareham, Hants England, PO16 0QB

Net: info@lasermoon.co.uk (mailto:info@lasermoon.co.uk)
support@lasermoon.co.uk (mailto:support@lasermoon.co.uk)

Provider's Description: We are specialists in UNIX/Linux Freeware and stock products from : Infomagic, PrimeTime Freeware, Yggdrasil, Linux Journal, SSC, O’Reilly Associates, Addison Wedsley, SWiM (Motif) 1.2.3 &1.2.4, Morse Telecommunications (Slackware Pro), VITAL (Crisp Editor), BAS-MARK (Language Products), Systek s.r.l (Application Builders), VERSASOFT (dBMAN), LINUX BIBLE and reference works. We are the sole UK/European/USA distributors for some of the above. If you don’t see what you want, please call us! We provide free technical support for the above products which are competitively priced. Dealers : Please contact us! We are looking for dealers in Europe and Aisa.

Availability:

CD-ROM, floppy disk and other media. ftp access for some products (i.e. SWiM)

Ordering:

We are pleased to accept : VISA, ACCESS, MASTERCARD, EuroCARD, Euro Cheque/Cheque/money order drawn on a UK bank in Pounds Sterling, Cash - (in desperation) Pounds Sterling (please send by recorded delivery).

Entry last modified:

14 Dec 1994.

7.4 Takelap Systems Ltd.

Distributor:

Takelap Systems Ltd. The Reddings, Court Robin Lane Llangwm, Usk, Gwent, United Kingdom NP5 1ET.

Net: info@ddrive.demon.co.uk (mailto:info@ddrive.demon.co.uk)

Vox: +44 1291 650357 Fax: +44 1291 650500 BBS: +44 1291 650567

Provider’s Description:

We supply the Infomagic Linux Developer’s Resource CD set. We also sell UNIX related books including ”Running Linux”.

Ordering:

UK price for the Infomagic CDs is gbp 22.33 including p&p and VAT. Credit cards accepted.

Miscellaneous:

Ask for our price list.

Entry last modified:

6 May 1995

8 Miscellany

This section lists distributors selling commercial software for Linux (such as Motif) and miscellany, such as documentation.

8.1 Unifix Custom Linux CD-ROM Mastering

Distributor:

Unifix Software GmbH, Braunschweig, Germany Postfach 4918 D-38039 Braunschweig, Germany

Vox: +49 (0)531 515161 Fax: +49 (0)531 515162

Provider’s Description:

We use our own (Linux based) premastering system to make custom CDs which include the customer’s programs in addition to our Linux distribution. This enables our customers to sell their Unix based applications together with operating system tested and ready to run. Our CD recorder makes it possible to create prototypes and very small series at interesting prices.

Availablility:

Ask us for more information or prices.

8.2 Linux Journal

Publisher:

Linux Journal P.O. Box 85867, Seattle, WA 98145-1867

Net: ljeditor@sunsite.unc.edu (mailto:ljeditor@sunsite.unc.edu)

Vox: (206) 782-7753 (206-PUBS-REF) Fax: 206-782-7191

Provider’s Description:

Linux Journal is a monthly publication covering the Linux Community. Most material in LJ is new (not reprinted from Usenet). Each issue includes columns and articles on Linux programming, Free Software Foundation issues, systems administration, Questions and Answers, interviews and more. LJ is a professional-quality magazine for the Linux community. Michael K. Johnson is the editor of Linux Journal.

Availability:

Subscriptions are $19/year (US), $24/year (Canada/Mexico), $29/year (elsewhere).

Ordering:

Payment can be made using VISA, MasterCard or American Express (make sure you include the credit card number, expiration date and signature). We can also accept checks in US $ on US banks. If you have a question concerning appropriate method of payment, phone or FAX our subscription numbers or send e-mail to subs@ssc.com For security reasons we discourage sending credit card numbers via e-mail.

Miscellaneous:

If you are interested in advertising in LJ, call, write or e-mail at linux@ssc.com. Article queries, new product announcements and other editorial material should be sent to our address above or e-mailed

to ljeditor@sunsite.unc.edu. Generic questions can be sent to linux@fylz.com.

Entry last modified:

11 Jan 1995.

8.3 Sequoia International Motif Development Package

Distributor:

Sequoia International, Inc. 600 West Hillsboro Blvd, Suite 300, Deerfield Beach, FL 33441,

Net: info@seq.com (mailto:info@seq.com)

Vox: (305) 480-6118 Fax: (305) 480-6198

Provider’s Description:

Sequoia International, Inc. has a complete Motif 1.2.3 Runtime and Development package called SWiM 1.2.3 available for $149.95. In addition to providing shared library versions of libXm and libMrm, the following is included in each package: The Window Manager (mwm), Shared Libraries (libXm, libMrm), (libXm, libMrm, libUil), UIL Compiler, Header and Include Files, Complete On-Line Manual Pages, Source Code to OSF/Motif Demo Programs, Complete OSF/Motif Users Guide.

Requirements:

Linux 0.99pl13 or higher, libc 4.4.4, XFree 2.0, 12M Free Disk, 8-12M RAM Suggested.

Ordering:

USA: Sequoia International, Inc. (305-480-6118), info@seq.com. Japan: Fortune Co., Ltd (03-5481-8974). England: Lasermoon Ltd. (+44-0-329-826444). Australia: Space Age Import-Export Proprietary (61-7-266-3418).

Entry last modified:

24 March 1994.

8.4 SSC Books and Linux Documentation Project Manuals

Distributor:

SSC, Inc. P.O. Box 55549 Seattle, WA 98155

Net: sales@ssc.com (mail:sales@ssc.com)

Vox: 206-782-7191 (206-PUBS-REF) Fax: 206-782-7191

Availability:

SSC has printed the Linux Installation and Getting Started manual, Version 2.2.1 to make it available for those who do not have the capability to print it themself and to support our customers who buy Linux distributions from us. The book is perfect-bound with a 2-color cover. SSC has also published the Network Administrator’s Guide from the LDP and will soon publish the Kernel Hacker’s Guide. In addition, SSC has published The Linux Sampler, a book which includes new material plus articles from back issues of Linux Journal. Other products include Linux T-shirts, mouse pads and bumper stickers.

Ordering:

Linux Installation and Getting Started, Version 2.1 is available for $12.95 plus shipping ($3 in the U.S.). The Linux Network Administrator’s Guide and The Linux Sampler are $18.95 each plus shipping. Finger bookshelf@ssc.com for information on these and other Linux books that we will make available. We can accept credit card orders (Visa, MasterCard or AmEx). Orders can be phoned in, faxed, mailed, or emailed.

Miscellaneous:

SSC also publishes a series of reference cards on Unix and Unix-related programs such as Emacs, VI, Korn Shell, C language, etc. SSC also sells the InfoMagic, Yggdrasil and Trans Ameritech distributions, Free Software Foundation books and a complete Linux package consisting of Linux Instllation and Getting Started, a Linux distribution on CD, 4 SSC Pocket References and a 1-year subscription to Linux Journal for $64.95. Call or e-mail sales@ssc.com for a free catalog.

Entry last modified:

11 Jan 1995.

8.5 Extent Verlag, LDP Distribution

Distributor:

Extent Verlag Berlin, Germany Postfach 12 66 48, D-10594 Berlin, Germany

Vox: +49 30 3244021 Fax: +49 30 3249685

Provider’s Description:

Extent has published the Linux Installation and Getting Started manual, Version 2.0 by Matt Welsh in order to make it available to Linux users within Germany and Europe. Inspite of being non-commercial, the manual was printed with 2540 dpi and bound in soft cover. Its handy paperback format is 148 x 210 mm. Other LDP manuals are being prepared as soon as possible. Extent also tries to make SLACKWARE available to everybody within Germany and Europe just for the ordinary diskette price. This means, you pay only the diskettes and get the newest SLACKWARE distribution for free, just in the sense of Free Software. The 3.5" diskettes are fully error-checked, of course.

Ordering:

Linux Installation and Getting Started, 192 pages. ISBN 3-926671-12-2. Costs within Germany DM 15.80 (tax and shipping included), within Europe DM 16.50 (shipping included). Full Slackware (latest version) on 50 diskettes. Within Germany, DM 100.00 plus DM 10.00 for shipping. Within Europe, DM 100.00 plus DM 20.00 for shipping. Ordering by sending cheque to address above or money order (don’t forget your address!) to Extent Verlag Berlin, Germany Postbank Berlin, BLZ 10010010, bank account 1769-104 or every book store.

Entry last modified:

25 March 1994.

8.6 UPython SLS and Slackware Diskette Labels

Distributor:

UPython Computer and Network Services

Provider’s Description:

Labels for Softlanding and slackware disks

Availability:

The labels that we print for you are available in several formats at

(sunsite.unc.edu:/pub/linux/distributions/SLSlabels).

Ordering:

For order information, send mail to DGray@uh.edu. Cost is $6.50 per set and domestic shipping in included. Overseas orders must send payment in US funds and add $1 for shipping. Texas residents add sales tax (6

Entry last modified:

11 Nov 1993.

9 Submissions To This Document

I encourage anyone and everyone who maintains a Linux distribution or mail order service (or consulting service, and the like), to submit information on their service to this HOWTO. It’s easy and fun, and it’s free advertising. This document is posted to many places and is archived (see the next section).

9.1 Types of submissions

We are interested in submissions for:

• Complete distributions of Linux software, available either via anonymous FTP, UUCP, or mail order. By ‘complete distribution’ we mean any set of software which can be used to build a complete Linux system from scratch.

• Layered products or individual software packages available only via mail order. If your software package is available via anonymous FTP, chances are people can find it. Software products only available via mail-order include commercial things such as Motif, and any other commercial software ported to Linux.

• Any other Linux-specific goods available via mail order, such as Linux-specific books and documentation, T-shirts, and assorted paraphrenalia. Note: our interest in the latter categories is marginal – this is a Linux Distributions HOWTO, not a catalogue of cutesy Linux junk. Whatever you’re selling will have to be (a) very useful, (b) very original, or (c) very funny, to get listed here.

If the number of submissions for services and layered products is large, I’ll create a separate HOWTO for these items.

9.2 How to submit

To submit an entry to this HOWTO, please send mail to esr@snark.thyrsus.com with the following information. This format is not machine-parsable; any of the fields may be any length that you wish, but I’d like to keep each entry down to, say, 50 lines.

Name:

Name of service or distribution

Distributor:

Name of company, person, etc. who distributes/maintains the service or distribution. Should include mail, email, phone contact information, and (if possible) a WWW reference.

Provider’s Description:

Description of the distribution or service that you provide. If this is a software distribution, please include information such as what software is included, versions, general overview of installation, requirements, and so on.

Internet Access:

Where your service or distribution is available over the Internet; typically an FTP address.

Ordering:

How to order your distribution or service, if applicable. Include prices, shipping information, methods of payment, etc.

Support:

Information on support terms and support contracts.

Miscellaneous:

Anything else that you find relevant.

Last Freeze Date:

Last freeze date of the current version(s). Also, your estimated update frequency.

9.3 Submission Guidelines¿

Please keep your entry as short as possible. If you need to include extensive information, please make a reference to where one can FTP or mail to get more information on your distribution; these entries are only meant to be pointers to where one can find information on your service or distribution.

If you provide more than one service or distribution, please use separate entries for each.

Some things (such as books, t-shirts, etc.) won’t fit exactly into this format; just be sure to include all relevant information. In other words, this ”entry form” is simply the bare minimum that you must include in your submission to the Distribution-HOWTO; feel free to change, add, or leave out ”fields” as you see fit.

I may edit your entries for conciseness and brevity, if I find any irrelevant information, or if the entry is overly verbose. Otherwise the content should remain the same.

When making submissions to the Distribution-HOWTO, you grant implicit permission for me to use the entries in other materials, such as books from the LDP, and other online documents. For example, information from the Distribution-HOWTO may be included in a published Linux book. If you do not want me to include your entry in materials other than the Distribution-HOWTO, please say so.